OpenStudy (anonymous):

express -5pi/6 radians in degrees in terms of pi help plz

OpenStudy (unklerhaukus):

use \[\pi^c=180°\]

OpenStudy (anonymous):

how would i do that

OpenStudy (unklerhaukus):

\[\frac{-5}6\pi^c=\frac{-5}6\times180°=\]

OpenStudy (anonymous):

-150 pi?

OpenStudy (unklerhaukus):

why is the pi still there?

OpenStudy (anonymous):

because we have to leave it in terms of pi

OpenStudy (unklerhaukus):

are you sure?

OpenStudy (unklerhaukus):

\[ -\frac{5\pi^c}6=-150°\]

OpenStudy (anonymous):

for the directions it says express radian meassures in terms of pi

OpenStudy (unklerhaukus):

yeah; the question dosent make much sense then

OpenStudy (anonymous):

i dnt know then

OpenStudy (anonymous):

express all radian measures in terms of pi express -5pi/6 radians in degrees?

OpenStudy (anonymous):

thank u for ure help :)